OpenStudy (anonymous):

Last one guys i promise ........the area of a rectangle is 112. The length is 7, and the width is √x. What is the value of x? What is the width of the rectangle? (Hint: The area of a rectangle is equal to length times width.)

ganeshie8 (ganeshie8):

use the hint

ganeshie8 (ganeshie8):

area = length * width 112 = 7 * \(\sqrt{x}\)

ganeshie8 (ganeshie8):

divide both sides by 7

ganeshie8 (ganeshie8):

112/7 = \(\sqrt{x}\) 16 = \(\sqrt{x}\) \(16^2 = x\) \(x = ? \)

OpenStudy (anonymous):

uhh, 5 ?